A balloon rubbed against denim gains a charge of -8.0C. What is the electric force between the balloon and the denim when the two are separated by a distance of 5.0 cm? (Assume that the charges are located at a point.)
1
Expert's answer
2020-02-25T10:22:17-0500

The quantity of electrostatic force between point charges is described by Coulomb's law.

"F=k\frac{q_1\cdot q_2}{r^2}" where "k=9\u00d710^9 N\u22c5m^2\u22c5C^{\u22122}" , "r=5.0cm=0.05m"

When electrified by friction, two bodies exchange charges. So there was a charge left on the denim +8.0C.

"F=9\cdot 10^9Nm^2C^{-2} \frac{-8.0C \cdot 8.0C}{(0.05)^2 m^2}=-2.3\cdot 10^{14}N"

Answer: The electric attractive force between the balloon and the denim is "2.3\cdot 10^{14}N"  

PS: It is possible that the value of the charge is incorrectly specified in the problem condition, which is why such a large force is obtained.